OpenStudy (anonymous):

Simplify: square root of 75

OpenStudy (anonymous):

\[\sqrt{75}=\sqrt{25*3}=5\sqrt{3}\]

OpenStudy (anonymous):

Work with the squareroot rules to solve the rest :) Try yourself, and then post your result and where you get stuck. Then we can continue to help.